boron has two isotopes 10b and 11b whose relative abundances are 20% and 80% then atomic wt. of boron is

Grade:12

2 Answers

Harendra Tiwari
6 years ago
Avg.atomic mass = €Ai x €Fi here €Ai is the sum of atomic mass and €fi is the sum of fractional distribution of isotope in nature since avg. atomic mass is the sum of the products of fractional aboundances of the isotope and their corresponding mass no. AVG atomic mass= .20x10+0.80x1=10.80
kunta gouthamreddy
6 years ago
Avg.atomic mass is obtained by multiplying the atomic mass of each isotope by its fractional abudance.
Ex.a and b are isotopes and with fractional abudance m and n respectively-
the Avg atomic mass=m*a+n*b/m+n(or)100.
Thus,Avg.atomic mass of boron=20*10+11*80/100=10.80